A curve on a flat road has curvature radius 111 m, and a caution sign urges drivers to slow to 60 km/h before negotiating the curve.Find the minimum frictional coefficient safe for driving at this speed along the curve of the road.
1
Expert's answer
2019-10-18T10:04:26-0400
"Ma= (mv^2)\/R"


"F_s=mgk"


"(mv^2)\/R=mgk"


"k= (mv^2)\/mgR"

"60 km\/h=600\/36 m\/s=50\/3 m\/s"

"k= v^2\/gR"


"k= \u301650\u3017^2\/(3^2\u22199,8\u2219111)\u22480,26"

Ansver: 0,26